### v2

# thanks: Equal contributionthanks: Equal contribution

Abstract
Run-and-tumble (RT) motion is commonly observed in flagellated microswimmers, arising from synchronous and asynchronous flagellar beating. In addition to hydrodynamic interactions, mechanical coupling has recently been recognized to play a key role in flagellar synchronization. To explore this, we design a macroscopic model system that comprises dry, self-propelled robots linked by a rigid rod to model a biflagellated microorganism. To mimic a low Reynolds number environment, we program each robot to undergo overdamped active Brownian (AB) motion. We find that such a system exhibits RT-like behavior, characterized by sharp tumbles and exponentially distributed run times, consistent with real microswimmers. We quantify tumbling frequency and demonstrate its tunability across experimental parameters. Additionally, we provide a theoretical model that reproduces our results, elucidating physical mechanisms governing RT dynamics.
pacs:
05.40.-a, 05.70.Ln, 45.70.Vn
preprint: APS/123-QED
Motility is one of the defining features of active and living organisms across length scales [1, 2, 3]. One prominent example is run-and-tumble (RT) motion commonly seen in swimming microorganisms living in a low Reynolds number environment [4, 5, 6, 7, 8]. Here, organisms trace relatively straight paths (runs) before abruptly changing to a new, randomly chosen direction (tumbles). It is known to originate from coordinated dynamics of multiple active units shaped as filamentous appendages, called flagella or cilia, decorating the body of microorganisms. These units exhibit intrinsic activity through their rhythmic beating, propelling the organism forward in a series of synchronous and asynchronous cycles [5, 9, 10, 11, 12, 13, 14, 15, 16].
Most organisms showing RT motion lack cognitive abilities or centralized control systems like the brain. Therefore, simple coupling rules between its active components must govern their dynamics. Since they swim in a fluid medium, attempts have been made to investigate the effects of hydrodynamic coupling between beating flagella on motility  [17, 18, 19, 20, 21, 22]. However, there is growing evidence that mechanical couplings between active components inside the organism’s body also play an important role. This is especially true for biflagellated organisms like Chlamydomonas reinhardtii [16, 23, 24, 25, 23, 26, 27]. Thus, it remains to be seen whether an artificial model system that demonstrates and validates the emergence of RT motion based on these mechanisms can be envisaged.
Figure 1: (a) Photograph of the experimental system featuring coupled robots. The pivot points, $P_{1}$ and $P_{2}$, allow the connecting rod to rotate freely in the horizontal plane. (b) Typical trajectories of free robots executing overdamped AB motion with $v_{a}$ = 5 cm s-1 for three different values of the rotational diffusion constant, $D_{r}$ (in rad2 s-1). (c) A schematic diagram highlighting key variables incorporated into the theoretical model. (d) A typical RT trajectory of the centroid point $C$ in the direction of black arrows, with $v_{a}$ = 5 cm s-1, $D_{r}$ = 0.06 rad2 s-1, $\delta$ = 3 cm, and $\alpha=90^{\circ}$. The color bar shows the value of $\beta$. (e) Zoomed-in views of typical tumble events under conditions of high (top) and low (bottom) substrate friction. We increased $v_{a}$ to 20 cm s-1 for low friction case to better illustrate the smoothening of the tumble event.
In the past, attempts have been made to mimic various dynamical features of microorganisms in synthetic systems [28, 29, 30, 31, 32, 33, 34]. Notwithstanding, an artificial analog system that accurately mimics RT motion with quick, straight run trajectories and slow, sharp tumbles is lacking. While there are reports of RT-like motion in vibrated granular particles [35] and self-propelling camphor boats [36], these studies often feature long tumble durations and very short runs. This behavior contrasts with microorganisms, where runs are much longer than tumbles [4, 5, 6]. This discrepancy likely originates due to finite inertial effects in these macroscale synthetic systems [37, 38, 39, 40], limiting their relevance to microscopic organisms that exclusively operate in inertialess limit. Therefore, for a macroscale artificial model system to accurately reproduce RT motion, it must similarly operate in an overdamped regime. To the best of our knowledge, no experimental system currently exists that fulfills these conditions while providing tunable statistical features of the RT motion. In this study, we introduce an experimental model system supported by a theoretical framework where a highly tunable RT-like motion spontaneously emerges in mechanically coupled self-propelled robots.
Our experimental system consists of two self-propelled robots connected with a rigid rod as shown in Fig. 1(a). The setup is conceived to model robots as active components and the rod as the body of an organism. Each robot is 7.5 cm in diameter and individually programmed to execute in-plane overdamped active Brownian (AB) motion along axes passing through their centers, aligned parallel to their wheels [41, 42] [also Supplemental Material (SM) Section IA]. This AB motion is characterized by a constant self-propulsion speed $v_{a}$ and the rotational diffusion constant $D_{r}$. Both robots are covered with 3D-printed caps featuring black, triangle-shaped markers, for tracking their in-plane positions and orientations. The AB trajectories of the single, non-connected robot are shown in Fig. 1(b) for three typical values of $D_{r}$ while keeping $v_{a}$ = 5 cm s-1 (see SM movie 1). Central to their overdamped feature is the fact that they propel forward using their wheels, which roll without slipping on a flat glass surface covered with a white sheet of paper. This provides sufficient traction for the wheels, making the effects of inertia negligible. The connecting rod of length $\ell=14.5$ cm is free to rotate about pivot points ($P_{1}$ and $P_{2}$) located on the off-centered, mirror-symmetric points on each robot [Figs. 1(a) &\&& (c)]. Here, $\delta$ is the distance of these points from robots’ centres, and $\alpha$ denotes the angle between the pivot-to-center line and the polarity axis. We vary $D_{r}$, $\delta$ and $\alpha$ between 0.06 - 1.33 rad2s-1, $1$ $3$ cm and $30^{\circ}$ $150^{\circ}$ respectively. $v_{a}$ is kept fixed at 5 cm s-1. A combination of self-propulsion and constraint force due to the rigid rod generates net torque around vertical axes passing through $P_{1}$ and $P_{2}$. These torques act stochastically in opposite handedness for the two robots indicated by curved arrows in Fig. 1(a). To summarize, this setup incorporates the fact that the run and tumble states are associated with translational and rotational motion which originate from an interplay between active forces and their corresponding moments, respectively. In what follows, we show that this simple-minded design is capable of showing RT-like motion.
We now develop a theoretical model for our system (for details see SM Secs. IIA and IIB). In our model, all dissipative forces, including frictional and resistive effects on a single robot with translational and rotational velocity $\mathbf{v}$\&& $\bm{\omega}$, are represented by a dissipative force and torque $\mathbf{F}_{\text{d}}=-\bm{\Gamma}\cdot\mathbf{v}$ and $\bm{\tau}_{\text{d}}=-\Gamma_{\tau}\bm{\omega}$, respectively. Here, $\bm{\Gamma}$ and $\Gamma_{\tau}$ denote the translational dissipation tensor and rotational dissipation coefficient, respectively. We assume that even when connected, robots remain overdamped. We also neglect the leading-order dependence of $\bm{\Gamma}$ and $\Gamma_{\tau}$ on $\mathbf{v}$ and $\bm{\omega}$. In experiments, the friction perpendicular to a robot’s polar axis (from wheel sliding) is significantly greater than the parallel friction (from rolling). Consequently, $\bm{\Gamma}$ adopts a non-scalar form dependent on the robot’s orientation $\mathbf{n}$, with dissipative forces given by $-\Gamma_{\parallel}\mathbf{v}_{\parallel}$ and $-\Gamma_{\perp}\mathbf{v}_{\perp}$, where $\mathbf{v}_{\parallel}$ and $\mathbf{v}_{\perp}$ are the velocity components along and perpendicular to $\mathbf{n}$, and $\Gamma_{\parallel},\Gamma_{\perp}>0$. Using these assumptions, we derive the equations of motion for the orientation $\theta$ of the unit vector $\mathbf{N}$, which is normal to the rod, the orientation angles $\theta_{i}$ of the robots ($i=1,2$) relative to $\mathbf{N}$, and the centroid, $C$, of the system, $\mathbf{r}=(\mathbf{r}_{1}+\mathbf{r}_{2})/2$, where $\mathbf{r}_{i}$ represents the center’s position of $i$th robot [see Fig. 1(c) and Eqs. (30–32) of SM]. All parameters match experimental values, except for two phenomenological ratios adjusted for optimal agreement: $\Gamma_{\parallel}/\Gamma_{\perp}=0.1$ and $14\Gamma_{\tau}/(\Gamma_{\parallel}d^{2})=$ 1, with robot diameter $d=7.5$ cm.
We start with parameters $\alpha=90^{\circ}$ and $\delta=3$ cm. When connected robots are set in motion, the point $C$ exhibits dynamics resembling fast, nearly straight trajectories (runs) with occurrences of abrupt sharp turns and sudden halts (tumbles). A typical trajectory is shown in Fig. 1(d) and SM movie 2 for $D_{r}=0.06$ rad2 s-1 and 1.33 rad2 s-1 in SM movie 3. We reproduce these dynamics in simulations by solving equations of $\theta_{i}$, $\theta$ and $\mathbf{r}_{i}$ of our theoretical model (SM movies 2 and 3). We highlight that it essential to minimize inertial time scales in capturing sudden, sharp tumbling events in experiments. This implies minimizing factors of $m/\Gamma_{\parallel}$, $m/\Gamma_{\perp}$ and $I/\Gamma_{\tau}$, where $m$ and $I$ are the mass and the moment of inertia of each robot respectively (Eqs. 4 and 5 in SM section IIA). To achieve the limit where these factors are non-negligible, we performed an experiment on a slippery glass surface coated with coconut oil. Interestingly, we find that a majority of tumbling events become significantly smoother (See Fig. 1(e) and SM movie 4), making them harder to distinguish from run states. In contrast, such smooth turns are rare on a frictional substrate. These findings are in accordance with previous studies in some active matter experiments where inertial effects are known to introduce time delays preventing abrupt changes in direction [37, 38, 39]. We further demonstrate that the RT motion does not occur when individual robots are programmed to perform Brownian dynamics (SM movie 5 and SM section IB), confirming its inherently active nature [43]. Henceforth, all results are obtained from experiments conducted on a frictional surface.
Figure 2: (a, b) Probability distribution of $\beta$ shows a pronounced peak near $\beta$ = 0. Insets: $\beta$ and $V$ are inversely related to each other for both experiment and simulation. Black dashed lines at $\beta$ = 60 sets a threshold to differentiate between run and tumble events. (c, d) Typical behaviour of $\beta$ as a function of time. $\tau_{t}$ and $\tau_{\text{run}}$ represent tumble duration and run-time respectively. (e, f) In both experiment and simulation, we observe exponentially decaying run times (dashed lines as a guide) and unimodal tumble durations exponentially. Experimental and simulation parameters are $v_{a}$ = 5 cm s-1, $D_{r}$ = 0.06 rad2 s-1, $\delta$ = 3 cm, and $\alpha=90^{\circ}$.
To carry out statistical analysis of the RT motion, we record long-time trajectories comprising hundreds of running and tumbling events for experiments and simulations. In experiments, we eliminate trajectories that experience disruptions due to boundary walls (see SM section IC for more details). We find that two key parameters effectively describe the RT motion: the pair angle $\beta(t)$ representing the instantaneous difference in the in-plane orientations of two robots [see schematic Fig. 1(c)] and the instantaneous speed $V(t)$ of the centroid $C$. A typical distribution of $\beta$ is shown in Figs. 2(a) and 2(b) for experiment and simulation, respectively. We also find that these parameters are inversely related to each other [insets of Figs. 2(a) and 2(b)] with a cluster of points at high $V$ and low $\beta$, which we identify as run states and vice versa for tumbles. To differentiate runs from tumbles quantitatively, we set an arbitrary threshold of $\beta=60^{\circ}$, corresponding to the half-maximum of the run peak. The typical behaviour of $\beta$ as a function of time is shown in Figs. 2(c) and 2(d) for experiment and simulation, respectively. By using $\beta=60^{\circ}$ as a reference line, we quantify tumbling duration ($\tau_{t}$) and run-time ($\tau_{\text{run}}$) from this time series for both experiment and simulation. The resulting distributions $\tau_{\text{run}}$ show exponential decay in each case [Figs. 2(e) and 2(f)]. Encouragingly, such exponential decays are ubiquitous in many microorganisms like swimming bacteria [4], Chlamydomonas Reinhartdii algae [5], and amoeba [6]. Our experiments also reveal that runs last much longer than tumbles [$<\tau_{\text{run}}>=7$ s, $<\tau_{t}>=3.5$ s for data presented in Fig. 2(c)] consistent with observations made in real biological systems [4, 5].
Figure 3: (a, b) The tumbling frequency, $\lambda$ increases with $D_{r}$ and $\delta$. (c, d) The phase diagram of $\lambda$ in $\alpha-D_{r}$ plane for $\delta$ = 3 cm. The region enclosed by the dashed line corresponds to $\lambda\approx 0$. (e, f) $\lambda$ exhibits a critical $D_{r}$ indicated by a black arrowhead beyond which tumbling emerges in the system. (g, h) The run speed, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ but decreases monotonically for $>90^{\circ}$. Here $D_{r}$ = 0.06 rad2 s-1 and error bars represent standard deviation.
Next, inspired by theoretical models used for modeling RT motion  [44, 45, 46, 47, 48, 49], we quantify this motion in terms of tumbling frequency, denoted by $\lambda$ which equals $<\tau_{\text{run}}^{-1}>$. The averaging is performed over all run events. We find that for $\alpha=90^{\circ}$, $\lambda$ shows a systematic variation as a function of $\delta$ and $D_{r}$, as shown in Fig. 3(a) and 3(b) for experiment and simulation, respectively. Our results show that both $D_{r}$ and $\delta$ promote tumbling in the system along the expected lines as they introduce stochasticity in the run state and increased torque, respectively [see Eq. (14)b of SM]. The system also shows an interesting phase behaviour when we vary the parameter $\alpha$. Experiments and simulation show that the run becomes considerably stable for $\alpha>90^{\circ}$. This is indicated by regions surrounded by dashed lines corresponding to vanishing $\lambda$ in Figs. 3(c) and (d) for experiment and simulation, respectively. When we plot $\lambda$ as a function of $D_{r}$ for various values of $\alpha$, we find a well-defined critical $D_{r}$ beyond which $\lambda$ increases abruptly indicated by black arrowheads in Figs. 3(e) and 3(f). On the contrary, for $\alpha\leq 90^{\circ}$, $\lambda$ shows a gradual increase with $D_{r}$. We further observe that the mean run speed, $V_{\text{run}}$, calculated by taking the average of $V$ while running, remains independent of $\alpha$ for $\alpha\leq 90^{\circ}$ and decreases with $\alpha$ as $\alpha$ exceeds $90^{\circ}$ [insets of Fig. 3(e) and (f)].
Now, we perform a theoretical analysis to explain the observed RT dynamics in our system and to provide a rationale for the $\alpha-D_{r}$ phase diagram. As observed in experiments and numerical simulations, the dynamics of the system at any given time are entirely governed by its internal configuration, which is represented by variables $\theta_{i}$ ($i$ = 1 and 2 for two robots) [Fig. 4(a)]. Therefore, we define another set of generalized angular coordinates: $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$ (see Fig. 4(a)). Note that $\beta$ equals the principal value of the angle $2\theta_{-}$ in the range $[-\pi,\pi]$ and $\theta_{+}$ indicates the average orientation of the robots with respect to the rod. Fig. 4(a) clearly shows that motility causes increasing and decreasing trends in $\theta_{1}$ and $\theta_{2}$, respectively. Therefore, by definition, $\theta_{-}$ also increases with time over large time scales. Moreover, for integers $n$, when $\theta_{-}=n\pi$, the robots are parallel to each other and thus move with the maximum possible speed of $v_{a}$. Similarly, when $\theta_{-}=(2n+1)\pi/2$, the system shows tumbling behavior. To elaborate further, we derive the equations of motion for $\theta_{\pm}$ from Eqs. (30) and (32) of SM, which are given as follows (also see SM section IIC):
$\dfrac{d\theta_{\pm}}{dt}=\mathcal{T}_{\pm}+\sqrt{2D_{\pm}}\eta_{\pm}(t),\$ (1)
where $\mathcal{T}_{+}=(2v_{a}/\ell)\sin\theta_{-}\left(\sin\theta_{+}-\ell\cos\theta_{+}\mathcal{G}\mathcal{H}\right)$, $\mathcal{T}_{-}=-(2\delta v_{a}/\Gamma_{\tau})\mathcal{G}\cos^{2}\theta_{+}\cos(\theta_{-}+\alpha)\sin\theta_{-}$, and $D_{\pm}$, $\mathcal{G}$ and $\mathcal{H}$ are the functions $\theta_{+}$ , $\theta_{-}$, and $\alpha$ (see SM section IIC). The strengths of the noise terms $D_{\pm}$ in Eq. (1) are of the order of $D_{r}$. $\eta_{\pm}(t)$ are the delta-correlated noise with zero mean and variance one.
Figure 4: (a) A schematic diagram clarifying the angular coordinates $(\theta_{+},\theta_{-})$. Here, $\theta_{\pm}=(\theta_{1}\pm\theta_{2})/2$. (b) Flow diagram of $(\theta_{+},\theta_{-})$ for $\alpha=120^{\circ}$ evaluated from the deterministic part $\mathcal{T}_{\pm}$ of Eq.  (1). $S_{1}$\&& $S_{2}$ and $U_{1}$\&& $U_{2}$ encircled by dotted lines represent stable and untable points respectively. Black dashed lines represent experimental trajectories for $D_{r}=0$. Schematic configurations of $S_{1}$ and $S_{2}$ are shown at the bottom. $\mathbf{n_{1}}$ and $\mathbf{n_{2}}$ are misaligned by an angle of $2\alpha-\pi=60^{\circ}$ in both cases. (c) Flow diagram for $\alpha=90^{\circ}$. Blue dotted lines at $\theta_{-}=n\pi$ are semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$, where $n$ denotes positive integers. Black dashed lines are the experimental trajectories of $D_{r}=0$. (d) Flow profile for $\alpha=60^{\circ}$. The system exhibits four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, and $U_{4}$). Solid blue lines at $\theta_{-}=n\pi$ represent configurations that are stable in $\theta_{-}$ and neutral in $\theta_{+}$. The black dashed experimental trajectories for the $D_{r}=0$ case. Three among infinitely many run configurations corresponding to $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ are shown at the bottom. The color bars in (b), (c) &\&& (d) correspond to the magnitude of the vector $(\mathcal{T}_{+},\mathcal{T}_{-})$ in rad s-1.
Let us now analyze the effect of the deterministic components, $\mathcal{T}_{\pm}$, on the dynamics of $\theta_{\pm}$. In Figs. 4(b)-4(d), we display the flow profiles of the angular coordinates $(\theta_{+},\theta_{-})$ in the absence of noise. The color map represents the magnitude $\mathcal{T}$ in rad s-1 of the vector with two components $(\mathcal{T}_{+},\mathcal{T}_{-})$. Note that the system is invariant under the transformations $\theta_{1}\to\theta_{1}\pm 2\pi$ and $\theta_{2}\to\theta_{2}\pm 2\pi$. Consequently, it remains invariant under $(\theta_{+},\theta_{-})\to(\theta_{+}\pm\pi,\theta_{-}\pm\pi)$ and $(\theta_{+},\theta_{-})\to(\theta_{+}\pm\pi,\theta_{-}\mp\pi)$. This implies that the first and third quadrants, as well as the second and fourth, represent the same set of systems. This symmetry is evident in the flow profiles as well.
We begin with the case where $\alpha>90^{\circ}$, with a typical flow profile shown for $\alpha=120^{\circ}$ in Fig. 4(b). Here, we find two stable ($S_{1}$, $S_{2}$) and two unstable ($U_{1}$, $U_{2}$) fixed points. We also observe an unstable fixed line along $\theta_{-}=n\pi$ which is less relevant to the results presented. Locations of stable points in the first and second quadrants are $(\theta_{+},\theta_{-})$ = $(\pi,3\pi/2-\alpha)$ and $(0,3\pi/2-\alpha)$ for $S_{1}$ and $S_{2}$ respectively, with system configurations shown below Fig. 4(b). Both configurations correspond to run states. Therefore, the system favors these stable points for low noise values, resulting in a pronounced run regime at low $D_{r}$ values, explaining the experimental results presented in Figs. 3(c) and (e). Interestingly, both $S_{1}$\&& $S_{2}$ correspond to non-aligned robots ($\mathbf{n_{1}}\nparallel\mathbf{n_{2}}$). To test this, we conduct experiments for the $D_{r}=0$ case using random initial conditions in $(\theta_{+},\theta_{-})$ and find that the system follows the predicted flow lines (black dashed lines in Fig. 4(b)) before eventually settling into one of the stable regions, indicated by the dashed circular regions (see SM movie 6). Moreover, as evident from configurations $S_{1}$&\&& $S_{2}$, the system’s run speed at the stable points is given by $v_{a}\sin\alpha$, which decreases with $\alpha$ for $\alpha>90^{\circ}$. This explains the decreasing trend in $V_{\text{run}}$ shown in insets of Figs. 3(e) and 3(f). We also observe that at low but finite $D_{r}$ values, the system runs corresponding to $\theta_{-}$ = $\pi/2-\alpha+n\pi$, which equals $\theta_{-}=n\pi-\pi/6$ for $\alpha=120^{\circ}$ while occasionally transitioning between $S_{1}$ and $S_{2}$ leading to tumbling events [See SM Section IID].
For $\alpha=90^{\circ}$, no stable fixed points are observed. Instead, there exists a fixed line [blue dashed line in Fig. 4(c)] at $\theta_{-}=n\pi$, which is semi-stable in $\theta_{-}$ and neutral in $\theta_{+}$ Therefore, as the system reaches this line, it runs for a while before sliding towards increasing $\theta_{-}$ due to noise. Experiments performed at $D_{r}=0$, in this case, show traced paths that agree well with theoretical values (see SM movie 7). Finally, for $\alpha<90^{\circ}$, and using $\alpha=60^{\circ}$ as a typical example, we find four unstable points ($U_{1}$, $U_{2}$, $U_{3}$, $U_{4}$) and a fixed line at $\theta_{-}=n\pi$ as shown in Fig. 4(d). This fixed line is neutral in $\theta_{+}$ while stable in $\theta_{-}$ corresponding to infinitely possible configurations with $\mathbf{n_{1}}\parallel\mathbf{n_{2}}$ with schematics of three typical configurations shown below the Fig. 4(d). Consequently, $V_{\text{run}}\approx v_{a}$ for $\alpha\leq 90^{\circ}$ as observed experimentally in inset of Fig. 3(e). Again, the grey dashed line indicates the experimental trajectory for $D_{r}=0$ (SM movie 8).
To conclude, inspired by microorganism motility, we present a robotic model system that replicates run-and-tumble-like (RT-like) motion. Our robots, driven by motorized wheels through mechanical gears and responding to microcontroller signals, undergo rolling-without-slipping motion. As a result, they faithfully emulate overdamped active motion, a hallmark of microorganisms operating at extremely low Reynolds numbers. We program both robots to exhibit an overdamped active Brownian (AB) motion. The rigid rod, attached to a pivot point on the robot’s body, rotates freely and facilitates rotational motion. By varying the pivot placement relative to the robot’s polarity axis, we observe a range of rich, tunable RT-like dynamics. We quantify the RT-like motion using the tumbling frequency and identify two key tuning parameters. The first, $\delta$, is the distance of the pivot from the robot’s centre, which promotes tumbling frequency by amplifying torque. The second, $\alpha$, is the angle between the pivot-to-center line and the polarity axis, which also adjusts the tumbling frequency based on rotational noise programmed inside each robot. The tunability of RT motion with these parameters might have implications in the critical role played by inter-flagella connection through the organism’s body in achieving synchronization cycles [23, 25]. We also developed a theoretical model that captures the essential features of our system and reproduces the experimental results, elucidating the intricate dependence of RT dynamics on $\delta$ and $\alpha$.
Acknowledgements - NK acknowledges financial support from DST-SERB for CRG grant number CRG/2020/002925 and IITB for the seed grant. HS acknowledges SERB for the SRG (grant no. SRG/2022/000061-G). SP thanks CSIR India for the research fellowship (File no: 09/087(1040)/2020-EMR-I).